OpenStudy (anonymous):

What is the equation of a line that has a slope of -5 and passes through the point (0, 2)?

hartnn (hartnn):

the equation of line in slope point form is y-y1 = m (x-x1) m= slope = -5 here (x1,y1) = point = (0,2) here now just put values. what u get ?

OpenStudy (anonymous):

idk...

hartnn (hartnn):

its just plugging in values y-y1 = m (x-x1) put y1 = 2, x1=0, m=-5 try it..

OpenStudy (anonymous):

kk thx i got it right

hartnn (hartnn):

ok, welcome ^_^